# Triadic Multi-party Voice Activity Projection for Turn-taking in Spoken Dialogue Systems

### v1

@@ 3.2 Model Comparison
We evaluated the VAP model by predicting the next speaker during periods of mutual silence. First, we located segments of mutual silence (i.e., no voice activity from any of the speakers) that were at least in duration. Next, we determined the next speaker. However, to prevent detecting backchannels or other short responses, we looked for speech that was at least in duration and made sure there was no other overlapping speech during that time. We evaluated the model’s expected probabilities from when the speaker begins up to in the past (i.e., the maximal amount the model predicts). Next, we aggregated the probabilities for the three speakers during this time. Finally, we compared the speaker with the highest probability during the mutual silence to the upcoming speaker to determine if the model’s prediction was accurate or not. We compared against a last speaker baseline where the model predicted the last speaker as the current speaker. Table 4 contains the accuracy for the three models. All models performed better than the respective baseline. The model trained only on spontaneous data had lower accuracy than those trained on attentive listening, as more participants speak simultaneously during discussions. As expected, performance accuracy was much higher for the attentive listening data, where a single speaker predominately leads the conversation.
@@
Figure 4 provides an example comparing the model’s predictions to ground truth labels during of dialogue. At , speaker 3 (green) begins a turn while briefly overlapping with speaker 2 (orange). The model indicates that speaker 3 has the highest output probability—correctly predicting the turn change—but underestimates speaker 2, illustrating the challenge of assigning probabilities during overlap, where predictions must exceed a threshold and are constrained to single-speaker outputs. Importantly, the model accurately predicts the switch from speaker 2 to speaker 3 even though both speakers overlap before and after the mutual silence.

### v2

@@ 3.2 Model Comparison
We evaluated the VAP model by predicting the next speaker during periods of mutual silence. First, we located segments of mutual silence (i.e., no voice activity from any of the speakers) that were at least in duration. Next, we determined the next speaker. However, to prevent detecting backchannels or other short responses, we looked for speech that was at least in duration and made sure there was no other overlapping speech during that time. We evaluated the model's expected probabilities from when the speaker begins up to in the past (i.e., the maximal amount the model predicts). Next, we aggregated the probabilities for the three speakers during this time. Finally, we compared the speaker with the highest probability during the mutual silence to the upcoming speaker to determine if the model's prediction was accurate or not. We compared against a last speaker baseline where the model predicted the last speaker as the current speaker. Table 4 contains the accuracy for the three models. All models performed better than the respective baseline. The model trained only on spontaneous data had lower accuracy than those trained on attentive listening, as more participants speak simultaneously during discussions. As expected, performance accuracy was much higher for the attentive listening data, where a single speaker predominately leads the conversation.
@@
Figure 4 provides an example comparing the model's predictions to ground truth labels during of dialogue. At , speaker 3 (green) begins a turn while briefly overlapping with speaker 2 (orange). The model indicates that speaker 3 has the highest output probability—correctly predicting the turn change—but underestimates speaker 2, illustrating the challenge of assigning probabilities during overlap, where predictions must exceed a threshold and are constrained to single-speaker outputs. Importantly, the model accurately predicts the switch from speaker 2 to speaker 3 even though both speakers overlap before and after the mutual silence.